OpenStudy (anonymous):

find f(6) if f(x)=x^2 divided by 3 plus x

OpenStudy (anonymous):

really could use some symbols here to make it clearer

OpenStudy (anonymous):

Just plug in 6 into the function and solve. f(3) = (3)^2 / (3+3) = 9 / 6

OpenStudy (sleepyjess):

I am thinking it is \(\sf f(x) = \dfrac{x^2}{3} + x\)

OpenStudy (anonymous):

If it's how sleepyjess says it is, then plug in 3 and compute it.

OpenStudy (sleepyjess):

You mean 6?

OpenStudy (anonymous):

Oh wow. 6 sorry.

OpenStudy (sleepyjess):

haha

OpenStudy (usukidoll):

how is this an integral?

OpenStudy (usukidoll):

a function is more like it

OpenStudy (sleepyjess):

integral? who said anything about an integral?

OpenStudy (usukidoll):

didn't the subject line says help me finish evaluating this integral

OpenStudy (sleepyjess):

\(\color{blue}{\text{Originally Posted by}}\) @I_need_math_help find f(6) if f(x)=x^2 divided by 3 plus x \(\color{blue}{\text{End of Quote}}\) No

OpenStudy (usukidoll):

look to the left of the screen

OpenStudy (anonymous):

This says nothing about an integral.

OpenStudy (usukidoll):

ugh I'll prove it hold up while I copy and paste it

OpenStudy (usukidoll):

johnnydicamillo Help me finish evaluating this integral

OpenStudy (sleepyjess):

that isn't this question....

OpenStudy (anonymous):

That's a different user.

OpenStudy (usukidoll):

oh man sorry about that XD!

OpenStudy (sleepyjess):

haha :)

OpenStudy (sleepyjess):

we all make mistakes :)

OpenStudy (usukidoll):

not my day...lol XD